OpenStudy (anonymous):

Identify the discontinuities, graph, and find the zeros of the rational function f(x) =(x^2-100)/(x+10)

OpenStudy (anonymous):

the discontinuity will occur when the denominator, (x+10)=0 factored form of f(x) is (x-10)(x+10)/(x+10) x+10=0 x=-10 The graph is discontinuous at x=-10. this is a removable, or point, discontinuity because (x+10) is a factor of the numerator. The zeros of the rational function occur when the numerator equals 0: (x+10)(x-10)=0 x=10, -10

OpenStudy (anonymous):

however, because the function is not defined at x=-10, there is only 1 zero for this function. that is x=10